# Selective Retrieval-Augmentation for Long-Tail Legal Text Classification

### v2

@@ Abstract
Legal text classification is a fundamental NLP task in the legal domain. Benchmark datasets in this area often exhibit a long-tail label distribution, where many labels are underrepresented, leading to poor model performance on rare classes. This paper proposes Selective Retrieval-Augmentation (SRA) as a solution to this problem. SRA focuses on augmenting samples belonging to low-frequency labels in the training set, preventing the introduction of noise for well-represented classes, and requires no changes to the model architecture. Retrieval is performed only from the training data to ensure there is no potential information leakage, removing the need for external corpora simultaneously. The proposed SRA method is tested on two legal text classification benchmark datasets with long-tail distributions: LEDGAR (single-label) and UNFAIR-ToS (multi-label). The results indicate that SRA attains higher micro-F1 and macro-F1 scores compared to all current LexGLUE baselines across both datasets, illustrating consistent improvements in long-tail legal text classification. The code repository is available at: https://github.com/Boheng-Mao/sra-legal
@@ 1 Introduction
This paper proposes Selective Retrieval-Augmentation (SRA), an effective approach designed for long-tail legal text classification. SRA focuses primarily on underrepresented labels. It retrieves semantically relevant contents entirely from the training set. Due to their length, semantic richness, and self-contained nature, legal documents are especially well-suited for in-domain retrieval without the reliance on external knowledge (Ariai et al., 2024). Augmentation is applied consistently during training, validation, and testing, and the classifier architecture remains unmodified.
@@
Propose Selective Retrieval-Augmentation (SRA), an approach that only retrieves less common labels to address the issue associated with class imbalance in supervised legal text classification.
@@
Utilize SRA alongside a RoBERTa-base model Liu et al. (2019) and perform extensive experiments on two legal benchmarks tailored for unique classification tasks: LEDGAR (single-label) and UNFAIR-ToS (multi-label). As illustrated in Figure 1, SRA performs better than all existing LexGLUE benchmark models on both datasets Chalkidis et al. (2022).
@@ 2.1 Legal Benchmarks and Long-tail Classification
A significant challenge in legal text classification is the long-tail distribution of label frequencies, exemplified by benchmarks such as LEDGAR, UNFAIR-ToS, and EUR-Lex Chalkidis et al. (2019). In the vast machine learning literature, methodologies such as logit adjustment (Menon et al., 2021) and decoupled representation learning (Kang et al., 2020) have been introduced, with recent surveys providing comprehensive analyses of long-tailed classification strategies (de Alvis and Seneviratne, 2024). Although these techniques alleviate imbalance at the loss or classifier level, they do not offer direct contextual assists for infrequent classes. The proposed SRA framework, on the other hand, adds to existing methods by selectively adding in-domain examples to labels that are not well represented.
@@ 2.2 Retrieval-Augmented Classification
Retrieval-augmented approaches were first proposed as a supplement to large language models in knowledge-intensive tasks by retrieving suitable information from external sources Lewis et al. (2020). The concept has been expanded into text classification, including frameworks created for multi-label text classification Chalkidis and Kementchedjhieva (2023), as well as approaches focusing on few-shot Yu et al. (2023) and zero-shot classifications Abdullahi et al. (2024). These frameworks demonstrated the flexibility of retrieval augmentation, but they rarely address the underrepresented labels in skewed datasets directly, and some are designed for open-domain cases where external information is essential. In contrast, SRA takes advantage of the standalone nature of legal texts, enabling retrieval directly within the training set to support the less common categories. This design enables SRA to provide augmentation without needing outside sources.
@@ 6 Conclusion
This paper presented Selective Retrieval-Augmentation (SRA) for the classification of legal texts, emphasizing the selective enhancement of low-frequency samples to mitigate noise associated with full augmentation. Results from tests on LEDGAR and UNFAIR-ToS showed noticeable improvements in both micro-F1 and macro-F1 scores. Analysis further proved that SRA is especially effective for rare categories. Future work involves the expansion of SRA to other long-tailed domains and its integration with more sophisticated retrieval and fine-tuning procedures.
@@
Selective Retrieval-Augmentation (SRA) consistently yields improvements. However, several limitations remain. First, the method utilizes heuristic rules (like a frequency-based cutoff) to decide which proportion to augment, and the best ratio may differ for other datasets. Second, the retrieval step is limited to fairly simple pipelines (TF-IDF + Sentence BERT), so the potential benefits of switching to more advanced retrieval and reranking techniques need further investigation. Third, since the experiments in this paper are limited to legal text classification, more experiments in other fields are required to verify whether the results are generalizable. Lastly, even with a relatively simple retrieval strategy, the augmentation process could make training less efficient, especially for larger datasets.

### v3

# Exploring Selective Retrieval-Augmentation for Long-Tail Legal Text Classification

@@ Abstract
Legal text classification is a fundamental NLP task in the legal domain. Benchmark datasets in this area often exhibit a long-tail label distribution, where many labels are underrepresented, leading to poor model performance on rare classes. This paper explores Selective Retrieval-Augmentation (SRA) as a proof-of-concept approach to this problem. SRA focuses on augmenting samples belonging to low-frequency labels in the training set, preventing the introduction of noise for well-represented classes, and requires no changes to the model architecture. Retrieval is performed only from the training data to ensure there is no potential information leakage, removing the need for external corpora simultaneously. SRA is tested on two legal text classification benchmark datasets with long-tail distributions: LEDGAR (single-label) and UNFAIR-ToS (multi-label). Results show that SRA achieves consistent gains in both micro-F1 and macro-F1 over LexGLUE baselines. The code repository is available at: https://github.com/Boheng-Mao/sra-legal
@@ 1 Introduction
This paper explores Selective Retrieval-Augmentation (SRA) for long-tail legal text classification. SRA focuses on underrepresented labels and retrieves semantically relevant content entirely from the training set. Due to their length, semantic richness, and self-contained nature, legal documents are particularly well-suited for in-domain retrieval without relying on external knowledge (Ariai et al., 2024). Augmentation is applied consistently during training, validation, and testing, while the classifier architecture remains unchanged. The evaluation setup deliberately assumes access to ground-truth labels to isolate the effect of selectively augmenting low-frequency classes. The limitations of this design are discussed later in the paper.
@@
Explore Selective Retrieval-Augmentation (SRA), a strategy that targets low-frequency classes for augmentation to reduce class imbalance in supervised legal text classification.
@@
Implement SRA with a RoBERTa-base model Liu et al. (2019) and evaluate it on two legal benchmarks: LEDGAR (single-label) and UNFAIR-ToS (multi-label). As illustrated in Figure 1, SRA achieves consistent gains over LexGLUE baselines across both datasets Chalkidis et al. (2022).
@@ 2.1 Legal Benchmarks and Long-tail Classification
A significant challenge in legal text classification is the long-tail distribution of label frequencies, exemplified by benchmarks such as LEDGAR, UNFAIR-ToS, and EUR-Lex Chalkidis et al. (2019). In the vast machine learning literature, methodologies such as logit adjustment (Menon et al., 2021) and decoupled representation learning (Kang et al., 2020) have been introduced, with recent surveys providing comprehensive analyses of long-tailed classification strategies (de Alvis and Seneviratne, 2024). Although these techniques alleviate imbalance at the loss or classifier level, they do not provide direct contextual support for infrequent classes. In contrast, Selective Retrieval-Augmentation (SRA) explores the idea of selectively incorporating in-domain examples for underrepresented labels, offering a complementary perspective to existing approaches.
@@ 2.2 Retrieval-Augmented Classification
Retrieval-augmented approaches were initially proposed as supplements to large language models in knowledge-intensive tasks by retrieving suitable information from external sources Lewis et al. (2020). The concept has since been extended to text classification, including frameworks developed for multi-label settings Chalkidis and Kementchedjhieva (2023), as well as approaches targeting few-shot Yu et al. (2023) and zero-shot classification Abdullahi et al. (2024). These studies have shown the flexibility of retrieval augmentation, but they rarely focus on underrepresented labels in skewed datasets, and many are designed for open-domain scenarios where external information is essential. Building on this line of work, Selective Retrieval-Augmentation (SRA) explores the use of in-domain retrieval within the training set to support less common categories, illustrating a way to provide augmentation without relying on outside sources.
@@ 6 Conclusion
This paper introduced Selective Retrieval-Augmentation (SRA) for long-tail legal text classification. The results show that directing augmentation toward tail classes can substantially improve both micro-F1 and macro-F1 compared with uniform augmentation and strong LexGLUE baselines. While the current implementation should be regarded as an exploratory study, the findings highlight the potential of tail-aware augmentation for addressing long-tail distributions in legal NLP.
@@
The selective retrieval-augmentation strategy uses ground-truth labels during validation and testing to determine whether a sample belongs to the low-frequency set defined from the training distribution. This assumption is unrealistic for deployment, where labels are unavailable at inference time. In this work, it is adopted only to provide a proof-of-concept, showing that tail-aware augmentation can be beneficial. Future work may investigate label-free alternatives, for example uncertainty-based criteria or auxiliary models that estimate whether an input is likely to come from a tail class.